A 100 kg elevator is descending with an acceleration od 1.0 ms-2. Calculate the tension in the suspending cable.
1
Expert's answer
2020-06-26T14:37:54-0400

Let's choose the upwards as the positive direction and apply the Newton's Second Law of Motion:


"T - mg = -ma,""T = mg - ma = m(g - a),""T = 100 \ kg \cdot (10 \ \dfrac{m}{s^2} - 1.0 \ \dfrac{m}{s^2}) = 900 \ N."

Answer:

"T = 900 \ N."